40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 26 19:48:09 2016, Lydia, Update, SUS, ITMX stuck again, ITMY whitening issue 44.png26.png
    Reply  Tue Sep 27 08:49:47 2016, Steve, Update, SUS, seismic activity is up 4.3mSaltonSee.pngITMXstuck.png
       Reply  Mon Oct 10 09:31:25 2016, Steve, Update, SUS, PRM damping restored 3.7mLomaLinda.pngETMY_UL_glitch.pngMC_glitcing?.png
       Reply  Tue Nov 1 09:03:08 2016, Steve, Update, SUS, seismic activity is up seismicActivity.png
          Reply  Thu Dec 8 07:56:05 2016, Steve, Update, PEM, EQ6.5m Northen CA eq6.5FerndaleCA.png16d_glitching-_trend.pngEQ6.5_&4.7mFerndaleCa.png
             Reply  Thu Dec 29 10:24:56 2016, Steve, Update, PEM, EQ5.7mHawthorn NV eq5.7HawthorneNV.png
    Reply  Tue Sep 27 18:04:50 2016, Lydia, Update, SUS, ITMX slow channels down, ITMY diagonalization update ITMY_diagsuccess.pdf
Message ID: 12518     Entry time: Mon Sep 26 19:48:09 2016     Reply to this: 12519   12520
Author: Lydia 
Type: Update 
Category: SUS 
Subject: ITMX stuck again, ITMY whitening issue 

This afternoon around 2:45, ITMX started ringing up at ~.9Hz for about a minute and then got stuck again. When I noticed this evening, I tried to free it with the alignment sliders but was unable to see any signal on UL or UR. It also looks like the damping for ITMY was turned off at the same time ITMX got stuck (not at the start of its ring up). SRM also has a spike in its motion at this time, and another one minute later that ended up with the LR OSEM at a much higher level, though the mirror does not appear to be stuck. We didn't see any strange behavior from any of the other optics.

Teng and I were working on diagnosing a problem with the ITMY UL whitening, but by the time we disconnected any applicable cables, the damping for ITMY was already off. Later we unplugged the ITMX PD whitening cables after verifying that the ITMX damping was also already off. This problem may have occured earlier, while Teng, Eric, and I were examining and pushing in the cables at 1X5 without unplugging anything.

We found that the reason for the bad phase on the ITMY free swing data is because the whitening filter for UL is not being properly turned on. We are in the process of investigating the source of this problem. Right now all the cables to the PD whitening boxes in 1X5 are switched between ITMY and ITMX.
 

Attachment 1: 44.png  61 kB  | Hide | Hide all
44.png
Attachment 2: 26.png  42 kB  | Hide | Hide all
26.png
ELOG V3.1.3-